The generic metal A forms an insoluble salt AB(s) and a complex AC5(aq).

The equilibrium concentrations in a solution of AC5 were found to be [A]=0.100 M[A]=0.100 M, [C]=0.0260 M[C]=0.0260 M, and [AC5]=0.100 M[AC5]=0.100 M. Determine the formation constant, Kf, of AC5.

?f=

The solubility of AB(s) in a 1.000 M solution of C(aq) is found to be 0.1310.131 M. What is the Ksp of AB?

?sp=
Answer #1

(x) Finding the formation constant - The possible reaction is- A +5c - Als concentration The formation constant ky is a hatio

Part B - Finding the Kep. The possible reaction occutang between Aband a is- AB (s) + 52 (aq) = Alglag) & Bear). To find the

0.017 16/ 4.887x10-3 - 305 1156 x 1073 10-3 - 3. 51156 The value of ksp = If 3,51156 8.416 * 109 (Ksp = 4.1724x1000)